The mass of a cubical box is 5 kgs and the dimensions are 30 x 30 x 30 cm. if the weight of the empty container is 1000kgs, calculate the weight of a container of size 4.5 x 6 x 3m filled with these boxes.
Please help me solve this problem!
Thank you

Answer by rothauserc(4718)

30 cm = 0.3 m, 0.3^3 = 0.027 m^3 is volume of a box
:
4.5 * 6 * 3 = 81 m^3 is volume of container
:
81 / 0.027 = 3000 boxes fit in container
:
3000 * 5 = 15000 kgs is the weight of 300 boxes
:
15000 + 1000 = 16000 kgs is the weight of the 300 boxes plus the weight of the container
: